OpenStudy (adi3):

Will medal Please help Find Two functions f and g such that: (a) f(g(x)) = (x+5)^3

OpenStudy (adi3):

@hartnn

OpenStudy (adi3):

@ganeshie8

OpenStudy (adi3):

Help pls

ganeshie8 (ganeshie8):

Try : f(x) = x^3 g(x) = x+5

OpenStudy (adi3):

thts right how did u do tht

OpenStudy (adi3):

@ganeshie8

OpenStudy (adi3):

@texaschic101

Nnesha (nnesha):

f(g(x)) means substitute x for g(x) function into f(x) \[\huge\rm f(\color{ReD}{g(x)})=(\color{ReD}{x+5})^3\] here is an example find f(g(x) ) when f(x) = x and g(x)= 2x so you should substitute x for 2x into f(x) function \[\large\rm f(\color{ReD}{g(x)})=\color{ReD}{x}~~~~~~~~~~~~~~f(\color{Red}{g(x)})=\color{Red}{2x}\]

Nnesha (nnesha):

make sense ? or do you need more explanation ?

OpenStudy (adi3):

i does not make sense

OpenStudy (adi3):

i don't get wht u tryna say

OpenStudy (adi3):

@ganeshie8

Nnesha (nnesha):

well f(g(x)) means f of `g` of x so we should replace x with the entire equation g(x)